Question 3

  • Time: 20 minutes
  • Total: 12 marks

Orville Company’s standard and actual costs per unit are provided below for the most recent period. During this time period 650 units were actually produced.

Standard

Actual

Materials:

     Standard: 2.5 metres at $2.7 per m.

$6.75

     Actual: 2.1 metres at $2.60 per m.

$5.46

Direct labour:

     Standard: 1.6 hrs. at $8.40 per hr.

13.44

     Actual: 1.7 hrs. at $8.50 per hr.

14.45

Variable overhead:

     Standard: 1.6 hrs. at $4.70 per hr.

7.52

     Actual: 1.7 hrs. at $4.20 per hr.

7.14

Total unit cost

$27.71

$27,05

For simplicity, assume there was no inventory of materials at the beginning or end of the period.

Required:

Given the information above, compute the following variances. Also indicate if the variances are favorable or unfavorable.

  1. Materials price variance                                                                         (2 marks)

  1. Materials quantity variance                                                                   (2 marks)
  1. Direct labour rate variance                                                                    (2 marks)
  1. Direct labour efficiency variance                                                          (2 marks)

  1. Variable overhead efficiency variance                                                 (2 marks)

                     6.        Variable overhead spending variance                                                                            (2 marks)
Answer #1

1. Materials price variance = AQ x (AP – SP) = 1365 x ($2.60 - $2.70) = $136.50 Favorable

AQ = 650 x 2.1 = 1365

2. Materials quantity variance = SP x (AQ – SQ) = $2.70 x (1365 – 1625) = $702 Favorable

SQ = 650 x 2.5 = 1625

3. Direct labour rate variance = AH x (AR – SR) = 1105 x ($8.50 - $8.40) = $110.50 Unfavorable

AH = 650 x 1.7 = 1105

4. Direct labour efficiency variance = SR x (AH – SH) = $8.40 x (1105 – 1040) = $546 Unfavorable

SH = 650 x 1.6 = 1040

5. Variable overhead efficiency variance = SR x (AH – SH) = $4.70 x (1105 – 1040) = $305.50 Unfavorable

6. Variable overhead spending variance = AH x (AR – SR) = 1105 x ($4.20 - $4.70) = $552.50 Favorable